# CropCraft: Inverse Procedural Modeling for 3D Reconstruction of Crop Plants

### v1

@@ Abstract
The ability to automatically build 3D digital twins of plants from images has countless applications in agriculture, environmental science, robotics, and other fields. However, current 3D reconstruction methods fail to recover complete shapes of plants due to heavy occlusion and complex geometries. In this work, we present a novel method for 3D reconstruction of agricultural crops based on optimizing a parametric model of plant morphology via inverse procedural modeling. Our method first estimates depth maps by fitting a neural radiance field and then employs Bayesian optimization to estimate plant morphological parameters that result in consistent depth renderings. The resulting 3D model is complete and biologically plausible. We validate our method on a dataset of real images of agricultural fields, and demonstrate that the reconstructions can be used for a variety of monitoring and simulation applications.
@@
In the following supplementary material, we provide additional experimental details (Sec. A) and additional experimental results (Sec. B). We invite readers to watch the supplementary video (supp.mp4) for further visualization.
@@ 1 Introduction
Plants are ubiquitous objects that appear all around the world and serve as the foundation for agriculture, which underpins our civilization’s growth and survival. The ability to automatically build 3D digital twins of plants from images has countless applications in agriculture, environmental science, robotics, and other fields. In particular, the development of such reconstruction methods in the context of agriculture will enable automatic, large-scale monitoring of crops. The collected data can provide decision support for farmers, aid carbon budgeting for decision-makers, support the development of new agricultural techniques, and inform the design of new genotypes [59, 10, 40]. All of these advances will contribute to increasing crop productivity, alleviating the rising food crisis of today’s world [4, 30].
However, 3D reconstruction of plants remains to be a challenging vision problem. Many plants, including most of those found in agriculture, are composed of complex arrangements of thin leaves and branches that heavily occlude one another. These low visibility conditions cause existing reconstruction pipelines to fail. Multi-view reconstruction methods such as those based on neural rendering [29, 46, 49, 54, 24] or multi-view stereo [53, 38, 50, 57] do not reconstruct the invisible regions of the scene, leading to incomplete plant shapes and extraneous geometry. Learning-based methods [15, 26] can overcome this issue but require large amounts of ground-truth 3D data for training, which is extremely difficult to collect for dense vegetation due to the same occlusion reasons.
@@
In this work, we present a novel method for 3D reconstruction of agricultural crops based on optimizing the parameters of a procedural plant morphology model. Our method combines the flexibility of data-driven neural reconstruction methods with the robust foundational knowledge in procedural models. To this end, we first use neural radiance field (NeRF) techniques [29, 46] to estimate the geometry of the visible surfaces in the scene. We then apply RANSAC [36] to estimate plant row locations and determine a canonical camera pose from which to render depth maps. Next, we render depth maps from both the NeRF and a virtual scene generated by a procedural generation model, and use Bayesian optimization to minimize a loss function with respect to the procedural model’s parameters.
The design of the loss function and the parameterization of the procedural model are crucial for obtaining a 3D reconstruction that is useful for field-level analysis applications. Importantly, we note that the fine-grained positions of each individual leaf are mostly irrelevant, as long as the aggregate shape characteristics of the crop canopy are faithful. Thus, we define our loss function on histogram statistics of the depth map and optimize a highly compact set of parameters to ensure that the key canopy characteristics for determining crop productivity are accurately captured without being distracted by fitting irrelevant details. And thanks to the strong priors imposed by the procedural model, this process jointly optimizes the complete plant shape, including portions not visible in the input images.
To validate our approach, we collect a multi-view dataset of real soybean and maize fields, paired with manual measurements of leaf area and leaf angle. We show that the proposed method can successfully reconstruct realistic crop canopies across different growth stages and estimate key canopy structure variables more accurately than baselines. We also show that the reconstructed 3D canopies can be directly fed into radiative transfer modeling software to provide accurate predictions of photosynthesis rates. The results highlight the potential for monitoring crop productivity directly from camera images instead of costly flux tower equipment. Our contributions are twofold:
@@
We introduce the first framework for image-based crop productivity quantification, making scalable carbon exchange monitoring a possibility.
@@ 2 Related Work
The task of reconstructing the 3D geometry of a scene given images is a longstanding problem in computer vision. Most reconstruction pipelines start from Structure-from-Motion (SfM), which infers camera parameters for the input images through joint optimization with detected keypoints [47, 37]. Afterwards, 3D geometry can be estimated based on photometric consistency across input views. Multi-view stereo (MVS) methods attempt to match correspondences across images and then triangulate the 3D coordinates of the points [53, 38, 50, 57]. Another approach for 3D reconstruction is based on fitting a neural radiance field (NeRF) to the scene [29, 46, 49, 54, 24, 25, 55]. NeRF is a high-fidelity and compact 3D scene representation that consists of a color field and a density field, both parameterized by neural networks and fitted to the input images via differentiable volume rendering [29]. After training, depth maps can be obtained by volume rendering with the density field. NeRF-based methods can model view-dependent effects and tend to give better surface geometry than MVS. Recently, an alternative scene representation based on 3D Gaussians has been proposed [22, 13], achieving higher visual fidelity and more efficient rendering than NeRF, but with less accurate geometry. In our approach, we use a standard off-the-shelf NeRF [46] to reconstruct visible surfaces of plants, which has been explored before [41, 2, 35, 16]. However, due to the inevitable occlusion in agricultural scenes, this is inadequate for full canopy reconstruction. To solve this issue, we turn to procedural generation models, which incorporate scientific knowledge and define a space of complete plant shapes that we can constrain our solution to.
@@
Inverse procedural modeling (IPM) refers to the problem of finding a procedure for generating a 3D representation of a given object or scene. The key design choices for IPM pipelines are the choice of procedural model and the method for optimizing the parameters of the model. Inverse procedural modeling has been applied for various input categories, with works focusing on building interiors [11, 5, 17, 28, 18], facades [8, 51], and driving scenes [20, 7]. The procedural models range from scene grammars [20, 12, 18] to constructive solid geometry trees [56] to L-systems [33, 27, 43, 23], which are especially popular for modeling botanical trees. To fit the model, a structural similarity measure is defined based on geometric and or semantic agreement. Depending on the type of model, which may require estimating procedural rules and or parameters [1], various search-based or dynamic programming-based optimization methods may be used to maximize the similarity measure. Unlike the multi-view reconstruction methods described above, inverse procedural modeling methods tend to focus less on photometrically exact reconstructions, aiming to obtain structurally similar approximations useful for simulation applications.
To the best of our knowledge, our work is the first to use inverse procedural modeling to reconstruct agricultural crops in the field. Some previous works have tackled individual plants [43, 23, 45], but modeling dense crop canopies poses a new set of challenges due to the high levels of occlusion and scale of the data. We propose an approach combining conventional multi-view reconstruction techniques with procedural generation to estimate 3D meshes that can be used for plant phenotyping, visualization, and simulation of biophysical processes.
@@ 3 Inverse Procedural Modeling for Crops
Our proposed method takes as input a set of images of a field of crops, and outputs a set of parameters that can be fed into a procedural generation model to produce a 3D mesh of the crops in the images. An overview of our method is provided in Fig. 2. By constraining the space of possible 3D shapes to the output space of the procedural model, we can ensure that the reconstructed plants are complete and have realistic leaf shapes and branch topology. Of course, we also need our reconstruction to match what we can observe in the input images. We achieve this by searching for parameters that minimize a loss function based on depth maps estimated from the input images and corresponding depth maps rendered from the reconstruction.
@@ 3.1 Procedural Morphology Model
Our procedural generation models are heavily based on existing work in the crop morphology literature. For soybean, we build upon the mCanopy model by Song et al. [42], and for maize, we build upon the coupled maize model by Qian et al. [34]. The mCanopy model consists of nodes of trifoliate leaves attached via a petiole to either the main plant stem or a branch stem. The maize model consists of large curved leaves growing from a single main stem. We introduce our own parameterization for each model, as illustrated in Fig. 3. The parameterization is designed to be flexible enough to capture the primary shape variations across instances of each species while remaining as low-dimensional as possible (5 parameters for soybean, 4 for maize). Note that the procedural generation process is in general stochastic; e.g. the model samples leaf angles according to a probability distribution. We refer readers to the supplementary for additional details about each model.
3.2 Depth Rendering
@@
The procedural generation model defines a space of 3D shapes that can represent plants of various stages of maturity, growing conditions, and cultivars. In order to search within this space for the shape that best matches the plants in input observations, we define a loss function based on depth maps estimated from the input images and corresponding depth maps rendered using the proposed shape.
To estimate depth maps from the input images, we use an off-the-shelf NeRF (Nerfacto [46]) and render the expected depth via volume rendering. In order to produce comparable depth maps for the procedurally generated scene, we need to know where to place the plants relative to the camera. Since it is extremely difficult to determine the precise base position of every individual plant under heavy occlusion, we assume that the plants are planted in rows with a known planting density and estimate the row locations using a heuristic approach that operates on a 3D point cloud extracted from the NeRF.
@@
Specifically, we render the depth for each input view and unproject each pixel to a 3D point. Then, we randomly sample a fixed number of points within a bounding box for the scene and perform voxel downsampling. Next, we segment the points into ground and plant using a threshold on their color in LAB space. We fit a plane to the ground points using RANSAC [36] and take an upper slice of the plant points based on their distance to the ground plane. Finally, we fit a line to each row by sequentially running RANSAC and removing the inliers for each fitted line. Additional details for the row-fitting procedure are provided in the supplementary.
@@
Given the ground plane and rows, we define a standardized camera to render from at a predetermined height above the center of the row with the most inliers, facing straight downwards with its -axis aligned with the row. This camera is used to render both the observation depth map from the NeRF and the depth map induced by the procedurally generated scene. We also use a color threshold and a bounding box to acquire foreground (plant) masks and corresponding to each depth map.
@@
Net Photosynthesis | Early Morning (6 AM) | Noon (12 PM) | Late Afternooon (4 PM) |
@@ 3.3 Loss Function
One of the key motivations for estimating the 3D structure of agricultural crops is to enable photosynthesis simulations that provide accurate predictions of primary productivity and other ecological variables. In general, we are interested in these variables at a large scale, at the level of patches or fields, not at the level of individual plants. Thus, the reconstruction of individual leaf locations may not be necessary as long as certain statistics of the aggregated crop canopy are accurate. With this in mind, we design a loss function around histogram statistics of the rendered depth maps, aiming to recover the key shape properties necessary for photosynthesis simulation while avoiding overfitting to irrelevant aspects of the scene geometry.
@@
To capture more information about surface normals, we include a loss on (histograms of) magnitudes of Sobel derivatives of the depth maps. We sum absolute values of the Sobel derivatives in both directions with a kernel size of 3.
@@ 3.4 Bayesian Optimization
Since the procedural generation model directly adds new mesh faces to create plants with different topology, the transformation from parameters to generated shape is not differentiable. Thus, it is difficult to minimize the loss function with respect to the parameters using gradient-based optimization methods. Instead, we employ Bayesian optimization, a black-box optimization method that is commonly used for hyperparameter tuning [39]. Bayesian optimization creates a surrogate for the objective function using Gaussian process regression, and then decides where to sample next by optimizing an acquisition function that combines uncertainty and the expected objective value [9]. We perform Bayesian optimization with a Matern kernel and the simple expected-improvement (EI) acquisition function [9] to estimate the procedural generation parameters that minimize our loss function. We find that, in many scenes, there may be multiple solutions with similar loss values. In order to make our method more robust to different random initialization seeds, we run the optimization 10 times (in parallel) and average the solutions.
@@ 4 Experiments
To validate our approach, we collected a dataset of multi-view images in real crop fields in the U.S. Midwest, paired with manual measurements of key morphological variables. For soybean, we collected at 3 geographical locations at 6 different time points throughout the growing season, for a total of 18 scenes. The images were captured covering a area using the Polycam app on an iPad Pro, around 50 per scene. The paired manual measurements consist of leaf areas and leaf angles (angle between surface normal and zenith direction). Leaf area was measured using a LAI-2200 scanner, and leaf angle was measured using a protractor. For maize, we collected 5 different time points in one location, with each scene containing around 500 images from a DJI Mavic UAV flying at a height of . The images were aligned using Agisoft Metashape. Here, only leaf area was measured, not leaf angles. Both the soybean and maize data will be made publicly available.
@@
Our evaluation metrics (and field measurements) are centered around the commonly considered canopy structure variables of leaf area index (LAI) and the leaf angle distribution. LAI is defined to be the total surface area of leaves per unit of ground area and is widely used in productivity models, climate models, and methods to estimate other vegetative surface properties [31]. Similarly, leaf angle is regarded as a key component of plant ecological strategy with significant impact on land surface properties such as carbon flux, surface temperature and spectral signature [52]. To assess how accurately our reconstructions capture these key traits, we consider the following metrics: LAI Error (LAIE) is the root mean squared error (RMSE) between the predicted and ground-truth (manually measured) LAI. LAI Percent Error (LAIPE) is the absolute percent error between the predicted and ground-truth LAI. Angle Mean Error (AME) is the RMSE between the mean of the predicted leaf angles and the mean of the ground-truth leaf angles. Angle Standard Deviation Error (ASDE) is the RMSE between the standard deviation of the predicted leaf angles and that of the ground-truth leaf angles.
@@
Ours | 0.69 | 0.15 | 12.07 | 7.39 | 0.97 | 0.26 |
Baselines.
We compare our reconstruction method with the following baselines: Poisson refers to applying Poisson surface reconstruction [21] on the point cloud extracted from NeRF, and then thresholding in 3D to remove the background geometry. All of the remaining mesh faces are then considered as leaf faces. MLP refers to using a learned multilayer perceptron (MLP) to predict procedural generation parameters given the histograms, mask area, and view height of the observation depth map. The MLP is trained on 20K pairs of histogram statistics and model parameters, synthesized by running the procedural generation model and rendering depth maps. Trust-Region is the same as our proposed method but uses a (constrained) trust-region method [6] for optimization instead of Bayesian optimization. This requires an estimate of the gradient of the loss function, which is obtained via 2-point finite difference estimation [44]. Random refers to uniformly randomly sampling values for the procedural generation parameters.
Implementation Details.
4.1 Canopy Reconstruction

### v2

# CropCraft: Complete Structural Characterization of Crop Plants From Images

@@ Abstract
The ability to automatically build 3D digital twins of plants from images has countless applications in agriculture, environmental science, robotics, and other fields. However, current 3D reconstruction methods fail to recover complete shapes of plants due to heavy occlusion and complex geometries. In this work, we present a novel method for 3D modeling of agricultural crops based on optimizing a parametric model of plant morphology via inverse procedural modeling. Our method first estimates depth maps by fitting a neural radiance field and then optimizes a specialized loss to estimate morphological parameters that result in consistent depth renderings. The resulting 3D model is complete and biologically plausible. We validate our method on a dataset of real images of agricultural fields, and demonstrate that the reconstructed canopies can be used for a variety of monitoring and simulation applications. Project page: https://ajzhai.github.io/CropCraft
@@
In the following supplementary material, we provide additional experimental details (Sec. F), additional experimental results (Sec. G), and a discussion of limitations (Sec. H). We invite readers to watch the supplementary video (supp.mp4) for further visualization.
@@ 1 Introduction
Plants are ubiquitous objects that appear all around the world and serve as the foundation for agriculture, which underpins our civilization’s growth and survival. The ability to automatically build 3D digital twins of plants from images has countless applications in agriculture, environmental science, robotics, and other fields. In particular, the development of such methods in the context of agriculture will enable automatic, large-scale monitoring of crops. The collected data can provide decision support for farmers, aid carbon budgeting for decision-makers, support the development of new agricultural techniques, and inform the design of new genotypes [68, 10, 49]. All of these advances will contribute to increasing crop productivity, alleviating the rising food crisis of today’s world [4, 37].
However, 3D reconstruction of plants remains to be a challenging vision problem. Many plants, including most of those found in agriculture, are composed of complex arrangements of thin leaves and branches that heavily occlude one another. These low visibility conditions cause existing reconstruction pipelines to fall short. Multi-view reconstruction methods such as those based on neural rendering [35, 54, 58, 63, 28] or multi-view stereo [62, 47, 59, 66] do not reconstruct the invisible regions of the scene, leading to incomplete plant shapes and extraneous geometry. Learning-based methods [17, 30, 13] can overcome this issue but require large amounts of ground-truth 3D data for training, which is extremely difficult to collect for dense vegetation due to the same occlusion reasons.
@@
In this work, we present a novel method for 3D modeling of agricultural crops based on optimizing the parameters of a procedural plant morphology model. Our method combines the flexibility of data-driven neural reconstruction methods with the robust foundational knowledge in procedural models. To this end, we first use neural radiance field (NeRF) techniques [35, 54] to estimate the geometry of the visible surfaces in the scene. We then apply heuristics to localize the planting rows and calculate a row-aligned camera pose from which to render depth maps. Next, we render depth maps from both the NeRF and a virtual scene generated by our procedural model, and use Bayesian optimization to minimize a specially designed loss function with respect to the procedural model’s parameters.
The design of the loss function and the parameterization of the procedural model are crucial for obtaining a 3D canopy model that is useful for field-level analysis applications. Importantly, for such applications, the fine-grained positions of each individual leaf are mostly irrelevant, as long as the aggregate shape characteristics of the crop canopy are faithful. We thus define our loss on histogram statistics of the depth map and optimize a highly compact set of parameters to ensure that the key canopy characteristics for determining crop productivity are accurately captured without being distracted by fitting irrelevant details. Thanks to the strong priors imposed by the procedural model, this process jointly optimizes the complete plant shape, including portions not visible in the input images.
To validate our approach, we collect a multi-view dataset of real soybean and maize fields, paired with manual measurements of leaf area and leaf angle. We show that the proposed method can successfully reconstruct realistic crop canopies across different growth stages and estimate key canopy structure variables more accurately than baselines. These variables are tightly linked to crop yield and are used extensively in crop science. We also show that the reconstructed 3D canopies can be directly fed into radiative transfer modeling software to provide accurate predictions of photosynthesis rates. The results highlight the potential for monitoring crop productivity directly from camera images instead of costly flux tower equipment. Our contributions are summarized below:
@@
We provide a framework for image-based growth and photosynthesis quantification, paving the way for scalable yield prediction and carbon exchange monitoring.
@@ 2 Related Work
The task of reconstructing the 3D geometry of a scene given images is a longstanding problem in computer vision. Most reconstruction pipelines start from Structure-from-Motion (SfM), which infers camera parameters for the input images through joint optimization with detected keypoints [55, 46]. Afterwards, 3D geometry can be estimated based on photometric consistency across input views. Multi-view stereo (MVS) methods attempt to match correspondences across images and then triangulate the 3D coordinates of the points [62, 47, 59, 66]. Another approach for 3D reconstruction is based on fitting a neural radiance field (NeRF) to the scene [35, 54, 58, 63, 28, 29, 64]. NeRF is a high-fidelity and compact 3D scene representation that consists of a color field and a density field, both parameterized by neural networks and fitted to the input images via differentiable volume rendering [35]. After training, depth maps can be obtained by volume rendering with the density field. NeRF-based methods can model view-dependent effects and tend to give better surface geometry than MVS. More recently, new methods based on 3D Gaussians have been proposed [25, 14, 19], achieving higher visual fidelity and efficiency than NeRF, but often with less accurate geometry. In our approach, we use an off-the-shelf NeRF [54] to reconstruct visible surfaces of plants, which has been explored before [50, 2, 44, 18]. However, due to the inevitable occlusion in agricultural scenes, this is inadequate for full canopy reconstruction. To solve this issue, we turn to procedural models, which incorporate scientific knowledge and define a space of complete plant shapes that we can constrain our solution to.
@@
Inverse procedural modeling (IPM) refers to the problem of finding a procedure for generating a 3D representation of a given object or scene. The key design choices for IPM pipelines are the choice of procedural model and the method for optimizing the parameters of the model. IPM has been applied for various input categories, with works focusing on building interiors [11, 5, 20, 33, 21], facades [34, 8, 60], driving scenes [23, 7], and forestry [52, 15, 36, 26, 27, 31]. The procedural models range from scene grammars [23, 12, 21] to constructive solid geometry trees [65] to L-systems [41, 38], which are especially popular for modeling botanical trees. To fit the model, a structural similarity measure is defined based on geometric and or semantic agreement. Depending on the type of model, which may require estimating procedural rules and or parameters [1], various search-based or dynamic programming-based optimization methods may be used to maximize the similarity measure. Unlike the reconstruction methods described above, inverse procedural modeling methods tend to focus less on photometrically exact reconstructions, aiming to obtain structurally similar approximations useful for simulation applications.
To the best of our knowledge, our work is the first to use IPM to reconstruct agricultural crops in the field and the first to integrate NeRFs into the IPM pipeline. Compared to individual plants [68, 31], modeling dense crop canopies poses a new set of challenges due to the high levels of occlusion. IPM methods for trees [52, 27] are inspirational but do not adequately constrain the plant topology for smaller crops. We propose a specialized IPM approach to estimate 3D crop meshes that can be used for plant phenotyping, visualization, and simulation of biophysical processes.
@@ 3 Inverse Procedural Modeling for Crops
Our proposed method takes as input a set of images of a field of crops, and outputs a set of parameters that can be fed into a procedural generation model to produce a 3D mesh of the crops in the images. An overview of our method is provided in Fig. 2. By constraining the space of possible 3D shapes to the output space of the procedural model, we can ensure that the reconstructed plants are complete and have realistic leaf shapes and branch topology. Of course, we also need our reconstruction to match what we can observe in the input images. We achieve this by searching for parameters that minimize a specialized loss function based on depth maps estimated from the input images and corresponding depth maps rendered from the reconstruction.
@@ 3.1 Procedural Morphology Model
Accurate crop modeling requires appropriate species-based shape constraints. Our procedural generation models are heavily based on existing work in the crop morphology literature. For soybean, we build upon the mCanopy model by Song et al. [51], and for maize, we build upon the coupled maize model by Qian et al. [42]. The mCanopy model consists of nodes of trifoliate leaves attached via a petiole to either the main plant stem or a branch stem. The maize model consists of large curved leaves growing from a single main stem. We introduce our own parameterization for each model, as illustrated in Fig. 3. The parameterization is designed to be flexible enough to capture the primary shape variations across instances of each species while remaining as low-dimensional as possible (5 parameters for soybean, 4 for maize). Note that the procedural generation process is in general stochastic; e.g. the model samples leaf angles according to a probability distribution. We refer readers to the supplementary for additional details about each model.
3.2 Pose Alignment for Depth Rendering
@@
The procedural generation model defines a space of 3D shapes that can represent plants of various stages of maturity, growing conditions, and cultivars. In order to search within this space for the shape that best matches the plants in input observations, we define a specialized loss function based on depth maps estimated from the input images and depth maps rendered using the proposed shape.
Note that, in order to properly compare depth maps between the real scene and the procedurally generated scene, the procedural plants should be placed in the same positions and orientations (relative to the camera) as in the real scene. However, it is practically impossible to determine the precise base position of every individual plant under heavy occlusion. Fortunately, it is also unnecessary to do so, since our applications demand characterizing the canopy overall, not each individual plant. Thus, we handle pose alignment by using geometric heuristics to estimate plant row locations, while carefully designing our loss function to be invariant to individual plant locations within each row.
@@
We assume that our crop plants are planted in rows with a known planting density (typically true). To estimate the row locations, we first use an off-the-shelf NeRF (Nerfacto [54]) to obtain a 3D point cloud of the visible surfaces. We render the depth for each input view and unproject each pixel to a 3D point. Then, we randomly sample a fixed number of points within a bounding box for the scene and perform voxel downsampling. Next, we segment the points into ground/plant using a threshold on their color in LAB space. We fit a plane to the ground points using RANSAC [45] and take an upper slice of the plant points based on their distance to the ground plane. Finally, we fit a line to each row by sequentially running RANSAC and removing the inliers for each fitted line. Additional details for the row-fitting procedure are provided in the supplementary.
@@
Given the ground plane and rows, we define a standardized camera to render from at a predetermined height above the center of the row with the most inliers, facing straight downwards with its -axis aligned with the row. This design will allow our loss to be invariant to the exact placement of plants along the -axis. The camera is used to render both the observation depth map from the NeRF and the depth map induced by the procedurally generated scene. Note that depth estimators such as VGGT [57] are unsuitable here since we are rendering depth from a novel (non-input) view. We also use a color threshold and a bounding box to acquire foreground (plant) masks and corresponding to each depth map.
@@ 3.3 Loss Function
One of the key motivations for estimating the 3D structure of agricultural crops is to enable photosynthesis simulations that provide accurate predictions of primary productivity and other ecological variables. Recall that we are interested in these variables at a large scale, at the level of patches or fields, not at the level of individual plants. Thus, the reconstruction of individual leaf locations is not necessary as long as certain statistics of the aggregated crop canopy are accurate. With this in mind, we design a loss function around histogram statistics of the rendered depth maps, aiming to recover the key shape properties necessary for photosynthesis simulation while avoiding overfitting to irrelevant aspects of the scene geometry.
@@
To capture more information about surface normals, we include another L2 loss on (histograms of) magnitudes of Sobel derivatives of the depth maps. We sum absolute values of the Sobel derivatives in both directions with a kernel size of 3.
@@ 3.4 Bayesian Optimization
Since the procedural generation model directly adds new mesh faces to create plants with different topology, the transformation from parameters to generated shape is not differentiable. Thus, it is difficult to minimize the loss function with respect to the parameters using gradient-based optimization methods. Instead, we employ Bayesian optimization, a black-box optimization method that is commonly used for hyperparameter tuning [48]. Bayesian optimization creates a surrogate for the objective function using Gaussian process regression, and then decides where to sample next by optimizing an acquisition function that combines uncertainty and the expected objective value [9]. We perform Bayesian optimization with a Matern kernel and the simple expected-improvement (EI) acquisition function [9] to estimate the procedural generation parameters that minimize our loss function. We find that, in many scenes, there may be multiple solutions with similar loss values. In order to make our method more robust to different random initialization seeds, we run the optimization 10 times (in parallel) and average the solutions, although this is not strictly necessary. An analysis of the distribution of the optimization results is provided in the supplementary.
@@ 4 Experiments
To validate our approach, we collected a dataset of multi-view images in real crop fields in the U.S. Midwest, paired with manual measurements of key morphological variables. For soybean, we collected at 3 geographical locations at 6 different time points throughout the growing season, for a total of 18 scenes. The images were captured covering a area using the Polycam app on an iPad Pro, around 50 per scene. The paired manual measurements consist of leaf areas and leaf angles (angle between surface normal and zenith direction). Leaf area was measured using a LAI-2200 scanner, and leaf angle was measured using a protractor. For maize, we collected 5 different time points in one location, with each scene containing around 500 images from a DJI Mavic UAV flying at a height of . The images were aligned using Agisoft Metashape. Here, only leaf area was measured, not leaf angles. Both the soybean and maize data are available through our project page.
@@
Our evaluation metrics (and field measurements) are centered around the commonly considered canopy structure variables of leaf area index (LAI) and the leaf angle distribution. LAI is defined to be the total surface area of leaves per unit of ground area and is widely used in productivity models, climate models, and methods to estimate other vegetative surface properties [39]. Similarly, leaf angle is regarded as a key component of plant ecological strategy with significant impact on land surface properties such as carbon flux, surface temperature and spectral signature [61]. To assess how accurately our reconstructions capture these key traits, we consider the following metrics: LAI Error (LAIE) is the root mean squared error (RMSE) between the predicted and ground-truth (manually measured) LAI. LAI Percent Error (LAIPE) is the absolute percent error between the predicted and ground-truth LAI. Angle Mean Error (AME) is the RMSE between the mean of the predicted leaf angles and the mean of the ground-truth leaf angles. Angle Standard Deviation Error (ASDE) is the RMSE between the standard deviation of the predicted leaf angles and that of the ground-truth leaf angles. We also evaluate the accuracy of the estimated stem structure and topology in the supplementary, using synthetic scenes instead of real scenes due to the difficulty of obtaining corresponding ground-truth measurements.
Baselines.
We compare our reconstruction method with a variety of baselines: Poisson refers to applying Poisson surface reconstruction [24] on the point cloud extracted from NeRF, and then thresholding in 3D to remove the background geometry. All of the remaining mesh faces are then considered as leaf faces. MLP refers to using a learned multilayer perceptron (MLP) to predict procedural generation parameters given the histograms, mask area, and view height of the observation depth map. The MLP is trained on 20K pairs of histogram statistics and model parameters, synthesized by running the procedural generation model and rendering depth maps. Trust-Region is the same as our proposed method but uses a (constrained) trust-region method [6] for optimization instead of Bayesian optimization. This requires an estimate of the gradient of the loss function, which is obtained via 2-point finite difference estimation [53]. Random refers to uniformly randomly sampling values for the procedural generation parameters.
Implementation Details.
@@
[…]
4.1 Canopy Shape Modeling